# Bayesian Signal Matching for Transfer Learning in ERP-Based Brain Computer Interface

### v1

@@ 5 Analysis of BCI data from Real Participants
We examined the performance of our BSM framework on real-participant data, where the real data were originally collected from the XXX Lab. A total of 41 participants were recruited to accomplish the experiment, but we restricted our study population to 9 participants with brain damages or neuro-degenerative diseases. Each participant completed both a calibration (training) session (TRN) and several free-typing sessions (FRT). For the calibration session, each participant was asked to copy a 19-character phrase “THE_QUICK_BROWN_FOX” including three spaces with 15 sequences, while wearing an EEG cap with 16 channels and sitting close to a monitor with a virtual keyboard. Details can be found Thompson et al., (2014).
The steps of real data analysis were as follows: First, we performed a series of data pre-processing techniques on the raw EEG signals of each participant. Next, we identified a proper source participant pool based on testing swLDA accuracy. We trained swLDA and SMGP using the calibration data from the new participant only as two reference methods and trained BSM-Mixture and MDWM including source data from source participants as two mixture models. We did not include other conventional ML methods because they did not have the borrowing version and we argued that swLDA could serve as the performance baseline for other ML methods. We demonstrated the result of a particular participant K151, a senior male diagnosed with amyotrophic lateral sclerosis (ALS) and showed the results of the remaining cohort in the supplementary materials.
@@ 5.1 Data Pre-processing Procedure and Model Fitting
We applied the band-pass filter between 0.5 Hz and 6 Hz and down-sampled the EEG signals with a decimation factor of 8. Next, we extracted a fixed response window of approximately 800 ms after each event such that the resulting EEG matrix contained 15 sequence segments, and each sequence contained 12 extracted event-related EEG signal segments. The resulting event-related EEG signal segment had around 800 ms, i.e., 25 sampling points per channel. Since the SNR of EEG signals is typically low, we further applied the xDAWN algorithm to enhance the evoked potentials (Rivet et al.,, 2009). The xDAWN method was an unsupervised algorithm that decomposed the raw EEG signals into relevant P300 ERPs and background noises and aimed to find a robust spatial filter that projected raw EEG signal segments onto the estimated P300 subspace by directly modeling noise in the objective function. We used all 16 channels as the signal input and selected the first two major components. We will denote raw EEG signal inputs and ERP functions after the xDAWN algorithm as transformed EEG signals and transformed ERP functions, respectively.
@@
To emphasize the calibration-less point, we used a small portion of K151’s training set and abundant training set from source participants for calibration. We identified 23 participants whose prediction accuracy of swLDA on FRT files exceeded 50% and used them as the source participant pool: For each source participant, we included the first ten sequence replications of the first five characters: “THE_Q” (17.5% of the entire calibration data), and the testing prediction accuracy was obtained with all FRT files (up to three sessions per participant). For K151, we included the first five sequence replications of the same five characters (8.8% of the entire calibration data). Therefore, the resulting EEG feature matrices for the new participant and source participants had sizes of (180, 2, 25) and (360, 2, 25), respectively. Here, 180, 360, 2, and 25 were the number of stimuli for new participant, the number of stimuli for source participants, the number of transformed ERP components and the length of EEG response window, respectively.
@@
The datasets were fitted with the equation (6). We applied two covariance kernels $\kappa_{1}$ and $\kappa_{0}$ for target and non-target ERP functions, respectively. Two kernels were characterized by $\gamma$-exponential kernels with length-scale and gamma hyper-parameters as $(0.2,1.2)$ and $(0.3,1.2)$, respectively. The pre-specified threshold $\delta_{Z}$ was 0.02. We ran the MCMC with three chains, with each chain containing 3000 burn-ins and 300 MCMC samples. The Gelman-Rubin statistics were computed to evaluate the convergence status of each model. We examined two perspectives on model fitting: matching and prediction. For matching, we reported the selection indicator $Z_{n}$. We also produced the transformed EEG signal function estimates by BSM and compared them to the results by BSM-Reference. For prediction, we reported the prediction accuracy on FRT files by BSM-Mixture, MDWM, swLDA, and SMGP up to five testing sequences. We only reported BSM-Mixture to utilize both advantages of BSM and BSM-Reference. We also included the Split-and-Merge Gaussian Process (SMGP) method (Ma et al.,, 2022) as another method for comparison. The SMGP method primarily focused on detecting the spatial-temporal discrepancies on brain activity between target and non-target ERP functions under the sequence-based analysis framework, which was different from our stimulus-based analysis framework.
5.2 Matching and Prediction Results
Figure 5 shows the inference results of K151. The first columns show the mean estimates and their 95% credible bands of the first two transformed ERP functions by BSM-Mixture. Only target ERP information was borrowed. The right two columns show the spatial pattern and spatial filter of the new participant’s partial training data by xDAWN. The average value of $\{Z_{n}\}$ among source participants was around 1.0%. Given the general low signal-to-noise ratio property, it indicated that the borrowing results were not dominated by certain source participant.
For Component 1, we observed a sinusoidal pattern such that the transformed ERP response achieved a minor positive peak around 50 ms post stimulus and a major negative peak around 200 ms post stimulus. The signal achieved its second positive peak around 350 ms post stimulus and gradually decreased to zero. Spatial pattern and filter plot indicated that the major negative peak was primarily attributed to parietal-occipital and occipital regions (channels PO7, PO8, and Oz). For Component 2, we observed a Mexican-hat pattern with a double-camel-hump feature. The transformed target ERP response achieved its first negative peak around 100 ms post stimulus and its two positive peaks around 250 ms and 450 ms post stimulus before it gradually decreased to zero. Spatial pattern and filter plot indicated that such double-positive peak was primarily attributed to central and parietal regions.
@@ 5.2 Matching and Prediction Results
Figure 5: The first column shows the mean estimates with 95% credible bands of the first two transformed ERP functions by BSM-Mixture. Only target ERP information was borrowed. The right two columns show the spatial patterns and spatial filters of the new participant’s partial training data by xDAWN.
@@
Table 4: Prediction accuracy (percentage) of all FRT files with up to five testing sequences by BSM-Mixture, MDWM, swLDA, and SMGP with five training sequences.
Testing Mixture Methods Reference Methods
@@
Table 4 shows the total prediction accuracy of all FRT files with respect to the testing sequences up to five by BSM-Mixture, MDWM, swLDA, and SMGP with five training sequences for K151. Overall, BSM-Mixture performed better than two reference methods. BSM-Mixture achieved 80% accuracy within 3 testing sequences, which was slightly faster than the existing method MDWM. Unfortunately, the SMGP method performed poorly in this case due to different assumptions in the analysis framework. The SMGP method treated one sequence of EEG data as an analysis unit, while our BSM framework treated one stimulus of EEG data as an analysis unit. Therefore, under the RCP design, the sample size for SMGP was 1/12 of that for BSM-Mixture, which led to inaccurate estimation of P300 response functions and poor prediction on FRT files.
In general, the advantage of BSM-Mixture was reflected in later testing sequences when participants were more likely to get distracted and have a higher heterogeneity level of EEG signals over time. Certain temporal intervals might become less separable and the attenuated amplitude estimates on these potentially confusing regions by BSM-Mixture actually reduced the negative effect of growing signal heterogeneity on prediction accuracy. For K151, temporal regions such as [500 ms, 800 ms] of Component 1, the double peaks of Component 2 by BSM-Mixture were attenuated than those by BSM-Reference. However, that the initial 100 ms and intervals around 350 ms of Component 1 became even more separable by BSM-Mixture might indicated that those intervals were robust over time and essential to maintain high prediction accuracy during FRT tasks.
Figure 6: A Spaghetti plot of character-level testing accuracy by BSM-Mixture, MDWM,, swLDA, and SMGP for 9 participants with at most 5 testing sequences and 5 training sequences. Means of each method were shown in red.
@@ 6 Discussion
In this article, we proposed a hierarchical Bayesian Signal Matching (BSM) framework to build a participant-semi-dependent, calibration-less framework with applications to P300 ERP-based Brain-Computer Interfaces. The BSM framework reduced the sample size for calibration of the new participant by borrowing data from pre-existing source participants’ pool at the participant level. The BSM framework specified the joint distribution of stimulus-related EEG signal inputs via a Bayesian hierarchical mixture model. Unlike conventional clustering approaches, our method specified the baseline cluster associated with the new participant and conducted pairwise comparisons between each source participant and the new participant by the criterion of log-likelihood. In this way, our method avoided the cluster label switching issue and could pre-compute model parameters from source participants to speed up the calibration in practice. In addition, since our method intrinsically borrowed the information from source participants via the binary selection indicators, we could use parameters from the baseline cluster to test directly without refitting the model with the augmented data. Finally, our hierarchical framework was base-model-free, such that it could be extended to any other classifiers with clear likelihood functions. We demonstrated the advantages of our method using extensive single- and multi-channel simulation studies and focused on the real data analysis among the cohort with brain injuries and neuro-degenerative diseases.
@@
In the real data analysis, we found that first transformed ERP response was associated with the effect of visual cortex (higher positive values within parietal-occipital and occipital regions, while close to zero elsewhere) and that the second transformed ERP response was associated with the typical P300 responses (higher positive values within central and parietal regions). The resulting pattern with double-camel humps might be due to the effect of latency jitter (Guy et al.,, 2021) or simply the superimposed effect of various signal contributions from central and parietal electrodes by xDAWN. The cross-participant results also suggested that major negative peaks before 200 ms and positive peaks around 350 ms post stimulus were preserved in the first two transformed ERP responses. The above scientific findings were consistent with the finding from our previous study Ma et al., (2022) that the performance of the P300 speller greatly depended on the effect of the visual cortex, in addition to electrodes located in central and parietal regions.
@@
In addition, compared to the simulation studies, the borrowing results of the real data analysis did not depend on certain source participants. Several reasons accounted for the results: First, the data generation mechanism of our method assumed that each observed EEG signal input was generated independently across stimuli, while the real data were extracted with overlapping EEG components between adjacent stimuli, and the overlapping temporal feature still existed after applying the xDAWN filter. It was a potential model mis-specification to apply our method to the real data, which might make it difficult to find a very good match on the participant level. Second, our similarity criterion was based on calculating the log-likelihood of the target transformed ERP functions using the generative base classifier, which might be more sensitive to noise than discriminant classifiers. In simulation studies, the noises were clearly defined and well controlled, while in the real data analysis, the noises might not be normally distributed, and the heterogeneity level of data increased not only across participants but also within the same participant over time, and the growing variability of data tended to diminish the magnitudes of transformed ERPs. Fortunately, since BSM modeled the selection indicator in a soft-threshold manner, it had a flavor of Bayesian averaging that generally attenuated the magnitudes of peaks. Such diminishing differences between target and non-target transformed ERP function estimates actually contributed to the prediction accuracy by suppressing the separability of temporal intervals with higher uncertainty.
@@
For cross-participant analysis, the transformed ERP function estimates of BSM were consistently attenuated compared to BSM-Reference, and the patterns of first two components were similar across nine participants. Figure 6 demonstrates the prediction accuracy of our target population cohort by BSM-Mixture, MDWM, swLDA, and SMGP. BSM-Mixture has similar prediction accuracy to MDWM, and they performed better than two reference methods. The improvements in BSM-Mixture and MDWM confirmed that people with brain injuries or neuro-degenerative diseases could definitely benefit from data borrowing, especially because they were the population who needed this type of assistance the most for daily communication.
@@
There were some limitations to our study. First, the generative base classifier we used was still sensitive to background noise compared to conventional discriminant classifiers. It might overly extract information from source participants to overfit the model. Second, the generative classifier might require careful tuning of hyperparameters such as the kernel hyperparameter selection of Gaussian process priors. The current strategy is to determine the values with heuristics, which might potentially affect the inference results and the prediction accuracy. Finally, although selection at the participant level was an intuitive idea, the characteristic of heterogeneity over time within the same participant suggested that we should have secondary sequence-level binary selection indicators. However, a side analysis of using sequence-level BSM-Mixture surprisingly performed worse than the one using participant-level selection indicators. It might be due to the fact that such refined selection indicators paired with generative base classifiers might overfit the training data and therefore had less power to generalize to the FRT data over time. Nevertheless, our main contribution was to demonstrate the advantages of the proposed hierarchical signal-matching framework, and the current framework could be expanded and improved in alternative ways.
@@
For future work, first, we would replace the generative base classifier with other robust discriminant classifiers, e.g., Riemannian geometry. The entire framework would remain unchanged except that we would specify the likelihood function by introducing Riemannian Gaussian distributions (Zanini et al.,, 2016), (Said et al.,, 2017). Even if we followed the generative pathway for classifiers, despite the poor performance of the original SMGP, it could be further improved by a slight modification. Since the original SMGP method focused on inferring on spatial-temporal brain activity under the P300 ERP-BCI setting rather than prediction, it usually required a large sample size for model fitting. Therefore, if we continued using the current generative Bayesian classifier, we would adopt the idea of split-and-merge to explicitly control the discrepancies between target and non-target (transformed) EEG signals, i.e., the temporal interval close to 800 ms on the stimulus level. Finally, we would incorporate our framework into calibration-free methods (Verhoeven et al.,, 2017). In addition to borrowing the training data from source participants and testing separately, we would directly start testing on the new participant and dynamically update the parameters of the baseline group as the data arrive, known as calibration-free methods.

### v2

@@ 5 Analysis of BCI data from Real Participants
We examine the performance of our BSM framework on real-participant data, where the real data were originally collected from the University of Michigan Direct Brain Interface Laboratory (UMDBI). A total of 41 participants were recruited to accomplish the experiment, and we demonstrated the results of nine participants with brain damages or neuro-degenerative diseases. Each participant completed one calibration (training) session (TRN) and up to three free-typing sessions (FRT). For the calibration session, each participant was asked to copy a 19-character phrase “THE_QUICK_BROWN_FOX” including three spaces with 15 sequences, while wearing an EEG cap with 16 channels and sitting close to a monitor with a virtual keyboard with a $6\times 6$ grid. Details of the experimental setup can be found Thompson et al., (2014).
The steps of the real data analysis are outlined as follows: First, we performed pre-processing techniques on raw EEG signals. Next, we identified a proper source participant pool based on testing swLDA accuracy. We trained swLDA and SMGP as two reference methods using the calibration data from the new participant only and trained BSM-Mixture and MDWM as two methods with data borrowing. The swLDA method was an existing efficient classifier. The MDWM method simply identified common features between the new participant and each source participant and did not perform explicit participant selection. Note that the SMGP method primarily focused on detecting the spatial-temporal discrepancies on brain activity between target and non-target ERP functions under the sequence-based analysis framework, while the other three methods treated stimulus-specific EEG data as feature vectors for classification. We did not include other conventional ML methods because they did not have the borrowing version. We demonstrated the detailed results of Participant K151, a senior male diagnosed with amyotrophic lateral sclerosis (ALS), and showed the results of remaining participants in the supplementary materials.
@@ 5.1 Data Pre-processing Procedure and Model Fitting
We applied the band-pass filter between 0.5 Hz and 6 Hz and down-sampled the EEG signals with a decimation factor of 8. Next, we extracted a fixed response window of approximately 800 ms after each event such that the resulting EEG matrix contained 15 sequence segments, and each sequence contained 12 extracted event-related EEG signal segments. The resulting event-related EEG signal segment had around 800 ms, i.e., 25 sampling points per channel. Since the SNR of EEG signals is typically low, we further applied the xDAWN filter to enhance the evoked potentials (Rivet et al.,, 2009). The xDAWN method was an unsupervised algorithm that decomposed the raw EEG signals into relevant P300 ERPs and background noises and aimed to find a robust spatial filter that projected raw EEG signal segments onto the estimated P300 subspace by directly modeling noise in the objective function. We used all 16 channels as the signal input and selected the first two major components. We will denote raw EEG signal inputs and ERP functions after the xDAWN algorithm as transformed EEG signals and transformed ERP functions, respectively.
@@
To emphasize the calibration-less point, we used a small portion of K151’s training set and abundant training set from source participants. We identified 23 participants whose prediction accuracy of swLDA on FRT files exceeded 50% and used them as the source participant pool: For each source participant, we included the first 10 sequence replications of the first five characters: “THE_Q”, and the testing prediction accuracy was obtained with all FRT files (up to three sessions per participant). For K151, we included the first five sequence replications of the same five characters. Therefore, the resulting EEG feature matrices for the new participant and source participants had sizes of (300, 2, 25) and (600, 2, 25), respectively. Here, 300, 600, 2, and 25 were the number of stimuli for new participant, the number of stimuli for source participants, the number of transformed ERP components and the length of EEG response window, respectively.
@@
The datasets were fitted with the equation (5). We applied two covariance kernels $\kappa_{1}$ and $\kappa_{0}$ for target and non-target ERP functions, respectively. Two kernels were characterized by $\gamma$-exponential kernels with length-scale and gamma hyper-parameters as $(0.2,1.2)$ and $(0.3,1.2)$, respectively. We ran the MCMC with three chains, with each chain containing 8,000 burn-ins and 1,000 MCMC samples. The Gelman-Rubin statistics were computed to evaluate the convergence status of each model. For matching, we reported the selection indicator $Z_{n}$. Due to the data heterogeneity, we expect to see partial matching. The largest posterior mean of $\{Z_{n}\}$ is around 0.2, the $\delta_{Z}$ should not exceed 0.2. In addition, most posterior means are around 4%, and we expect to see two or three source participants to match with the new participant. We set the $\delta_{Z}=0.1$, i.e., the mean of $2/24$ and $3/24$. We also produced the transformed EEG signal function estimates by BSM and compared them to the results by BSM-Reference. For prediction, we reported the prediction accuracy on FRT files by BSM-Mixture, MDWM, swLDA, and SMGP up to five testing sequences. We only reported BSM-Mixture to utilize both advantages of BSM and BSM-Reference.
5.2 Inference and Prediction Results
@@
For the estimation of mean-level transformed signals, Figure 5 shows the inference results of K151. The first columns show the mean estimates and their 95% credible bands of the first two transformed ERP functions by BSM-Mixture. Only target ERP information was borrowed. The middle and right columns show the spatial pattern and spatial filter of the new participant’s partial training data by xDAWN, respectively.
Figure 5: The first column shows the mean estimates with 95% credible bands of the first two transformed ERP functions by BSM-Mixture. Only target ERP information is borrowed. The middle and right columns show the spatial patterns and spatial filters of the new participant’s partial training data by xDAWN, respectively.
For Component 1, we observed a sinusoidal pattern such that the transformed ERP response achieved a minor positive peak around 50 ms post stimulus and a major negative peak around 200 ms post stimulus. The signal achieved its second positive peak around 350 ms post stimulus and gradually decreased to zero. Spatial pattern and filter plots indicated that the major negative peak was primarily attributed to parietal-occipital and occipital regions (channels PO7, PO8, and Oz). For Component 2, we observed a Mexican-hat pattern with a double-camel-hump feature. The transformed target ERP response achieved its first negative peak around 100 ms post stimulus and its two positive peaks around 250 ms and 450 ms post stimulus before it gradually decreased to zero. Spatial pattern and filter plots indicated that such double-positive peak might be associated with the signals in central and parietal regions. The credible bands of BSM-Mixture were generally wider than BSM-Reference due to the uncertainty of pairwise comparisons among source participants.
For matching results, based on the P300 ERP estimation from BSM-Reference, we expect one or two source participants to match with the new participant, so we set the prior probability of $Z_{n}=1$ as $\Pr\{Z_{n}=1\}=2/24\approx 8\%$. And we set the $\delta_{Z}$ to be slightly higher than the prior probability, i.e., 10%. Among 23 source participants, five source participants had posterior probability of being merged with K151 greater than or equal to 10%. Of all, K111 showed the greatest similarity to K151, with a posterior probability of 25%, followed by K158 (16%), K183 (15%), and K172 (13%). The mean and median values of posterior probability $\Pr\{Z_{n}=1\}$ were 6.4% and 4.0%, respectively. Details of the matching results can be found in the Section S5.2 in the Supplementary material. Given the general low signal-to-noise ratio in EEG data, our borrowing results based on mean-level separation effects were not dominated by any particular source participant.
@@
Table 2: Prediction accuracy (percentage) of all FRT files with up to six testing sequences by BSM-Mixture, MDWM, swLDA, and SMGP with five training sequences.
Testing Borrowing Methods Reference Methods
@@
6 94.6% 95.9% 83.8% 59.8%
Table 2 shows the total prediction accuracy of all FRT files with respect to the testing sequences up to six by BSM-Mixture, MDWM, swLDA, and SMGP with five training sequences for K151. Overall, BSM-Mixture performed better than two reference methods. BSM-Mixture achieves an 85.1% accuracy within 2 testing sequences, which is slightly faster than the existing method MDWM. Unfortunately, the SMGP method performs poorly in this case due to different assumptions in the analysis framework. The SMGP method treats one sequence of EEG data as an analysis unit, while our BSM framework treats one stimulus of EEG data as an analysis unit. Therefore, under the RCP design, the sample size for SMGP is 1/12 of that for BSM-Mixture, which leads to inaccurate estimation of P300 response functions and poor prediction on FRT files.
In general, the advantage of BSM-Mixture is reflected in later testing sequences when participants are more likely to get distracted and have a higher heterogeneity level of EEG signals over time. Certain temporal intervals may become less separable and the attenuated amplitude estimates on these potentially confusing regions by BSM-Mixture actually reduce the negative effect of growing signal heterogeneity on prediction accuracy. For K151, temporal regions such as [500 ms, 800 ms] of Component 1, the double peaks of Component 2 by BSM-Mixture are both attenuated than those by BSM-Reference. However, that the initial 100 ms and intervals around 350 ms of Component 1 becomes even more separable by BSM-Mixture may suggest that those intervals are robust over time and essential to maintain high prediction accuracy during FRT tasks.
@@ 6 Discussion
In this article, we propose a hierarchical Bayesian Signal Matching (BSM) framework to build a participant-semi-dependent, calibration-less framework with applications to P300 ERP-based Brain-Computer Interfaces. The BSM framework reduces the sample size for calibration of the new participant by borrowing data from pre-existing source participants’ pool at the participant level. The BSM framework specifies the joint distribution of stimulus-related EEG signal inputs via a Bayesian hierarchical mixture model. Unlike conventional clustering approaches, our method specifies the baseline cluster associated with the new participant and conducts pairwise comparisons between each source participant and the new participant by the criterion of log-likelihood. In this way, our method avoids the cluster label switching issue and can pre-compute model parameters from source participants to speed up the calibration. In addition, since our method intrinsically borrows the information from source participants via the binary selection indicators, we can use parameters from the baseline cluster to test directly without refitting the model with the augmented data. Finally, our hierarchical framework is base-model-free, such that it can be extended to any other classifiers with clear likelihood functions. We demonstrate the advantages of our method using extensive single- and multi-channel simulation studies and focus on the cohort with brain damages.
@@
In the real data analysis, we find that first transformed ERP response is associated with the effect of visual cortex (higher positive values within parietal-occipital and occipital regions, while close to zero elsewhere) and that the second transformed ERP response is associated with the typical P300 responses (higher positive values within central and parietal regions). The resulting pattern with double-camel humps may be due to the effect of latency jitter (Guy et al.,, 2021) or simply the superimposed effect of various signal contributions from central and parietal electrodes by xDAWN. The cross-participant results also suggest that major negative peaks before 200 ms and positive peaks around 350 ms post stimulus are preserved in the first two transformed ERP responses. The above scientific findings are consistent with the finding from our previous study Ma et al., (2022) that the performance of the P300 speller greatly depends on the effect of the visual cortex, in addition to electrodes located in central and parietal regions.
@@
In addition, compared to the simulation studies, the borrowing results of the real data analysis show partial matching. Several reasons accounted for the results: First, the data generation mechanism of our method assumed that each observed EEG signal input is generated independently across stimuli, while the real data were extracted with overlapping EEG components between adjacent stimuli, and the overlapping temporal features still exist after applying the xDAWN filter. It is a potential model mis-specification to apply our method to the real data, which might make it difficult to find a very good match on the participant level. Second, our similarity criterion is based on calculating the log-likelihood of the target transformed ERP functions using the generative base classifier, which may be more sensitive to noise than discriminant classifiers. In simulation studies, the noises are clearly defined and well controlled. In the real data analysis, the noises may not be normally distributed, and the heterogeneity level of data increases not only across participants but also within the same participant over time, and the growing variability of data tends to compromise the separation effects of transformed ERPs. Fortunately, since BSM models the selection indicator in a soft-threshold manner, it has a flavor of Bayesian averaging that generally attenuates the magnitudes of peaks. Such diminishing differences between target and non-target transformed ERP function estimates suppress the weights of “ambiguous” temporal intervals and reduce the negative impact of false positives.
Figure 6: A Spaghetti plot of character-level testing accuracy by BSM-Mixture, MDWM,, swLDA, and SMGP for 9 participants with five training sequences. Means of each method are shown in red. A minimum accuracy of 0.7 for practical usage is shown.
@@
For cross-participant analysis, although we acknowledge the difference among the large and diverse source population, the features of interests for task-based, P300 ERP speller system are more consistent and easier to identify, compared to resting-state EEG data. We argue that the similarity exists among the participants who use this application. The transformed ERP function estimates of BSM are consistently attenuated compared to BSM-Reference, and the patterns of first two components are similar across nine participants. Figure 6 demonstrates the prediction accuracy of our target population cohort by BSM-Mixture, MDWM, swLDA, and SMGP. BSM-Mixture has similar prediction accuracy to MDWM, and they perform better than two reference methods. The minimal useful accuracy for communication is considered to be 70% (Kübler and Neumann,, 2005). The improvement in BSM-Mixture shows that people with brain injuries or neuro-degenerative diseases benefit from data borrowing.
@@
There are some limitations to our study. First, the generative base classifier we used is still sensitive to background noise compared to other discriminant classifiers. It might overly extract information from source participants to overfit the model. Second, the generative classifier might require tuning hyper-parameters such as the kernel hyper-parameter selection of Gaussian process priors although our sensitivity analysis suggests that the inference and prediction are robust with respect to moderate changes of kernel hyper-parameters. The current strategy is to determine the values with heuristics, which might potentially affect the inference results and the prediction accuracy. Finally, although selection at the participant level is an intuitive idea, the characteristic of heterogeneity over time within the same participant suggested that we should have secondary sequence-level binary selection indicators. However, a side analysis of using sequence-level BSM-Mixture surprisingly performs worse than the one using participant-level selection indicators. It might be due to the fact that such refined selection indicators paired with generative base classifiers might overfit the training data and therefore has less power to generalize to the FRT data over time. Nevertheless, our main contribution is to demonstrate the advantages of the proposed hierarchical signal-matching framework, and the current framework could definitely be expanded and improved.
@@
For future work, first, we would replace the generative base classifier with other robust discriminant classifiers, e.g., Riemannian geometry. The entire framework would remain unchanged except that we would specify the likelihood function by introducing Riemannian Gaussian distributions (Zanini et al.,, 2016), (Said et al.,, 2017). Even if we followed the generative pathway for classifiers, despite the poor performance of the original SMGP, it could be further improved by a slight modification. Since the original SMGP method focused on inferring on spatial-temporal brain activity under the P300 ERP-BCI setting rather than prediction, it usually required a large sample size for model fitting. Therefore, if we continued using the current generative Bayesian classifier, we would adopt the idea of split-and-merge to explicitly control the discrepancies between target and non-target (transformed) EEG signals, i.e., the temporal interval close to 800 ms on the stimulus level. Finally, we would incorporate our framework into calibration-free methods (Verhoeven et al.,, 2017). In addition to borrowing the training data from source participants and testing separately, we would directly start testing on the new participant and dynamically update the parameters of the baseline group as the data arrive.